OpenStudy (anonymous):

Will fan and medal please help!!!!!!!!!!!! What is the solution set of 4x = y and 2x2 − y = 0? {(0.5, -0.5), (2, -8)} {(0, 0), (2, 8)} {(1.5, 2), (8, 2)} {(-1.5, -2), (2, 8)}

OpenStudy (anonymous):

@caozeyuan

OpenStudy (caozeyuan):

2x2?!

OpenStudy (caozeyuan):

2x^2 or something?

OpenStudy (anonymous):

Yea sorry it's 2x^2

OpenStudy (caozeyuan):

well, if 4x=y then 2x^2-y=2x^2-4x=x^2-2x=0

OpenStudy (anonymous):

Soooo B???

OpenStudy (caozeyuan):

yep